# A General Framework for Inference-time Scaling and Steering of Diffusion Models

### v3

@@ Abstract
Diffusion models produce impressive results in modalities ranging from images and video to protein design and text. However, generating samples with user-specified properties remains a challenge. Recent research proposes fine-tuning models to maximize rewards that capture desired properties, but these methods require expensive training and are prone to mode collapse. In this work, we propose Feynman-Kac (fk) steering, an inference-time framework for steering diffusion models with reward functions. fk steering works by sampling a system of multiple interacting diffusion processes, called particles, and resampling particles at intermediate steps based on scores computed using functions called potentials. Potentials are defined using rewards for intermediate states and are selected such that a high value indicates that the particle will yield a high-reward sample. We explore various choices of potentials, intermediate rewards, and samplers. We evaluate fk steering on text-to-image and text diffusion models. For steering text-to-image models with a human preference reward, we find that fk steering a 0.8B parameter model outperforms a 2.6B parameter fine-tuned model on prompt fidelity, with faster sampling and no training. For steering text diffusion models with rewards for text quality and specific text attributes, we find that fk steering generates lower perplexity, more linguistically acceptable outputs and enables gradient-free control of attributes like toxicity. Our results demonstrate that inference-time scaling and steering of diffusion models – even with off-the-shelf rewards – can provide significant sample quality gains and controllability benefits. Code is available here. **footnotetext: Denotes equal authorship. Correspondence to rsinghal@nyu.edu, zfh2000@columbia.edu, rst306@nyu.edu.
@@ 1 Introduction
We introduce Feynman-Kac steering (fk steering), an inference-time steering framework for sampling tilted distributions with diffusion models. fk steering enables guidance with arbitrary reward functions, differentiable or otherwise, for both discrete and continuous-state space models. fk steering makes use of a rare-event simulation method, fk interacting particle systems (Moral, 2004; Hairer and Weare, 2014; Vestal et al., 2008; Chopin et al., 2020). This technique enables the generation of samples with high-rewards, which may be rare events under the original model $p_{\theta}(\mathbf{x}_{0})$.
@@
  • Provides a valuable method for generating samples with (rare) specified attributes. As an example, the method can be used to generate toxic language to support efforts in responsible AI (Zhao et al., 2024), including for red-teaming model behavior, and for fine-turning and steering generalist language models to recognize and reject toxic language input. On toxicity, without gradient guidance, fk steering can increase the toxicity rate of a text diffusion model from $0.3\%$ to $69.7\%$ with $k=8$ particles, outperforming both gradient guidance and best-of-$n$.
@@ 2 Related Work
fk steering builds on top of recent work such as tds (Wu et al., 2023b) and others (Trippe et al., 2022; Cardoso et al., 2023; Dou and Song, 2024) that propose inference-time techniques, which make use of particle-based sampling methods for sampling from the unnormalized target distributions.
@@
The sampling procedure for traditional autoregressive language models poses a challenge for smc approaches, since smc typically requires estimating the reward of the full sequence given its prefix. Lew et al. (2023) address this challenge by limiting to rewards calculated within a fixed look-ahead window. In contrast, Zhao et al. (2024) learn intermediate twisting potentials to marginalize over the remaining elements of a particular partial sequence. In our work, we demonstrate that intermediate estimates from diffusion models, for attributes like toxicity, can be used even with off-the-shelf reward models by evaluating the reward models on intermediate denoised estimates (see fig. 5).
@@ 4 Experiments
For fk steering, we resample $50$ times during the inference process (every $10$ steps for SSD-LM and every $20$ or MDLM). For all models we use $\lambda=10.0$, and return the highest scoring sample at $t=0$. Following Han et al. (2023), we generate $20$ continuations of length $50$ using their $15$ controllable generation prompts. In addition to fk steering, we evaluate base model performance and best-of-$n$. As a baseline to compare quality improvements from scaling inference steps (Gong et al., 2023; Sahoo et al., 2024), we also include results for SSD-LM with $T=5000$ (versus $T=500$). We additionally include results for GPT2-Medium (Radford et al., 2019b) as an auto-regressive baseline. We evaluate all methods using perplexity computed with GPT2-XL (Radford et al., 2019b), CoLA acceptability, and distinct uni/bi/trigrams (Dist-1/2/3) (Han et al., 2023; Li et al., 2016). Additonal details on our text experiments are included in B.
@@
GPT2-medium $1$ 13.8 85.3 53/90/94
SSD-LM $1$ 21.9 67.7 46/82/91
SSD-LMT×10 $1$ 18.9 72.3 45/81/90
SSD-LM + FK(GPT2) $4$ 11.3 85.0 40/73/85
SSD-LM + FK(Trigram) $4$ 13.9 79.0 40/76/88
SSD-LM + FK(CoLA) $4$ 18.0 97.7 46/81/91
SSD-LM + BoN(GPT2) $4$ 12.9 77.3 40/73/85
SSD-LM + BoN(Trigram) $4$ 15.9 75.0 42/77/88
SSD-LM + BoN(CoLA) $4$ 19.8 94.7 46/82/91
SSD-LM + BoN(GPT2) $8$ 11.3 81.3 40/73/85
SSD-LM + BoN(Trigram) $8$ 14.4 74.7 42/76/88
SSD-LM + BoN(CoLA) $8$ 18.6 96.0 45/81/91
MDLM $1$ 79.2 30.0 56/91/94
MDLM + FK(GPT2) $4$ 51.5 43.0 51/85/92
MDLM + FK(Trigram) $4$ 37.2 35.3 49/85/92
MDLM + FK(CoLA) $4$ 69.6 68.3 58/88/91
MDLM + BoN(GPT2) $4$ 57.1 38.0 55/89/93
MDLM + BoN(Trigram) $4$ 51.3 33.3 53/89/94
MDLM + BoN(CoLA) $4$ 68.6 61.0 57/91/94
MDLM + BoN(GPT2) $8$ 46.2 37.7 53/86/92
MDLM + BoN(Trigram) $8$ 44.7 35.0 52/88/93
MDLM + BoN(CoLA) $8$ 64.6 73.3 57/91/93
Table 3: Text sample quality results metrics. We sample texts of length $50$ from all models and score perplexity with GPT2-XL, CoLA acceptability, and Dist-1/2/3 diversity metrics. Both SSD-LM and GPT-medium have $355$ million parameters. MDLM is a smaller model with $170$ million parameters.
@@ 4.3 Studying different choices of rewards in fk steering
SSD-LM 1.0% 1.33% 21.9 46/82/91
SSD-LM + Gradient-Guidance 18.0% 20.0% 39.5 52/88/93
MDLM 0.3% 2.3% 79.2 56/91/94
Gradient-Free (SSD-LM)
SSD-LM + BoN(4) 1.7% 4.3% 22.5 46/81/91
SSD-LM + BoN(8) 5.3% 10.7% 22.6 46/82/91
SSD-LM + FK($k=4$) 8.0% 14.0% 22.1 46/82/91
SSD-LM + FK($k=4$, learned $r_{\phi}$) 14.0% 21.3% 26.6 45/83/92
SSD-LM + FK($k=8$) 27.0% 27.3% 23.8 46/83/92
SSD-LM + FK($k=8$, learned $r_{\phi}$) 39.0% 41.0% 26.2 45/83/91
Gradient-Free (MDLM)
MDLM + BoN(4) 2.0% 4.0% 83.2 56/89/93
MDLM + BoN(8) 3.7% 10.3% 89.5 58/91/93
MDLM + FK($k=4$) 22.0% 28.3% 72.6 55/88/92
MDLM + FK($k=4$, many $r_{\phi}$) 47.0% 47.7% 78.8 57/89/92
MDLM + FK($k=8$) 52.3% 43.3% 77.8 56/88/91
MDLM + FK($k=8$, many $r_{\phi}$) 69.7% 57.3% 65.7 56/87/90
Table 5: Toxicity results. We evaluate the toxicity of the generated samples with both the classifier used for steering the model and a separate holdout classifier. We report GPT2-XL perplexity and distinct $n$-gram metrics.
@@
We consider the task of red-teaming toxicity, a rare attribute identified in only $1\%$ of base SSD-LM samples and $0.3\%$ of MDLM samples. Toxicity presents an undesirable behavior for language models. Here, we examine whether fk steering enables examining rare but dangerous model behavior, a critical factor considered before deploying systems (Zhao et al., 2024). In this experiment, we run fk steering with the base text diffusion models with SSD-LM (Han et al., 2023) and MDLM models. We use the base model as the proposal generator. As a baseline, we compare against gradient guidance for SSD-LM. For reward, we use a popular toxicity classifier (Logacheva et al., 2022). [^9] https://huggingface.co/s-nlp/roberta_toxicity_classifier
@@
  • Learned rewards improves controllability: With $8$ particles, fk steering SSD-LM outperforms gradient guidance on holdout toxicity accuracy ($27.3\%$ vs $20.0\%$), and dramatically improves on perplexity ($23.8$ vs $39.5$). Using improved intermediate rewards, learned from real data, improves toxicity accuracy even further to $41\%$.
  • Using many-sample $r_{\phi}$ improves controllability: fk steering MDLM with $k=8$ achieves an accuracy of $47.7\%$. fk steering outperforms best-of-$n$ sampling with both $4,8$ particles. Using improved intermediate rewards for MDLM more samples for intermediate rewards improves performance even further to $57.3\%$ ($69.7\%$ on the guidance reward model).
@@
In the previous section, we observed that for the same number of particles, using either learned rewards or multiple samples improves performance. For instance, using $k=8$ with SSDLM without learning the intermediate rewards gets a toxicity rate of $27\%$, higher than the learned reward’s toxicity rate of $14\%$; using the learned rewards improves the toxicity rate to $39\%$. However, these rewards come with an added computational cost, either in training or more evaluations of the reward model, presenting a trade-off in terms of using extra computational resources either for more particles or for better intermediate rewards.
@@ 5 Conclusion
Promising future directions include exploring the value of varying the numbers of particles at inference time. For applications like protein design, dynamically assigning greater numbers of particles to promising regions could enable generating a large number of diverse candidates. Additionally, we believe it is crucial to better understand the limits of inference-time particle scaling and the corresponding compute-performance trade-offs.

### v4

@@ Abstract
Diffusion models produce impressive results in modalities ranging from images and video to protein design and text. However, generating samples with user-specified properties remains a challenge. Recent research proposes fine-tuning models to maximize rewards that capture desired properties, but these methods require expensive training and are prone to mode collapse. In this work, we present Feynman-Kac (fk) steering, an inference-time framework for steering diffusion models with reward functions. fk steering works by sampling a system of multiple interacting diffusion processes, called particles, and resampling particles at intermediate steps based on scores computed using functions called potentials. Potentials are defined using rewards for intermediate states and are selected such that a high value indicates that the particle will yield a high-reward sample. We explore various choices of potentials, intermediate rewards, and samplers. We evaluate fk steering on text-to-image and text diffusion models. For steering text-to-image models with a human preference reward, we find that fk steering a 0.8B parameter model outperforms a 2.6B parameter fine-tuned model on prompt fidelity, with faster sampling and no training. For steering text diffusion models with rewards for text quality and specific text attributes, we find that fk steering generates lower perplexity, more linguistically acceptable outputs and enables gradient-free control of attributes like toxicity. Our results demonstrate that inference-time scaling and steering of diffusion models – even with off-the-shelf rewards – can provide significant sample quality gains and controllability benefits. Code is available here. **footnotetext: Denotes equal authorship. Correspondence to rsinghal@nyu.edu, zfh2000@columbia.edu, rst306@nyu.edu.
@@ 1 Introduction
In this work, we present Feynman-Kac steering (fk steering), a flexible framework for steering diffusion-based generative models with arbitrary rewards that uses fk interacting particle system methods (Moral, 2004; Vestal et al., 2008). We generalize previous works that define Feynman-Kac measures to conditionally sample diffusion models (Trippe et al., 2022; Wu et al., 2023b; Chung et al., 2022; Janati et al., 2024). fk steering enables guidance with arbitrary reward functions, differentiable or otherwise, for both discrete and continuous-state models. The approach makes use of a rare-event simulation method, Feynman-Kac interacting particle system (fk-ips) (Moral, 2004; Del Moral and Garnier, 2005; Hairer and Weare, 2014; Vestal et al., 2008). fk-ips enables the generation of samples with high-rewards, which may be rare events under the original model $p_{\theta}(\mathbf{x})$.
@@
  • Provides a valuable method for generating samples with (rare) specified attributes. As an example, the method can be used to generate toxic language to support efforts in responsible AI (Zhao et al., 2024a), including for red-teaming model behavior, and for fine-turning and steering generalist language models to recognize and reject toxic language input. On toxicity, without gradient guidance, fk steering can increase the toxicity rate of a text diffusion model from $0.3\%$ to $64.7\%$ with $k=8$ particles, and outperforms both gradient guidance and best-of-$n$.
@@ 2 Related Work
fk steering builds on top of recent works that sample from Feynman-Kac path distributions for conditional sampling with diffusion models, either using particle-based sampling (Trippe et al., 2022; Wu et al., 2023b; Cardoso et al., 2023; Dou and Song, 2024; Zhao et al., 2024b) or gradient-based sampling (Chung et al., 2022; Janati et al., 2024).
@@
The sampling procedure for traditional autoregressive language models poses a challenge for smc approaches, since smc typically requires estimating the reward of the full sequence given its prefix. Lew et al. (2023) address this challenge by limiting to rewards calculated within a fixed look-ahead window. In contrast, Zhao et al. (2024a) learn intermediate twisting potentials to marginalize over the remaining elements of a particular partial sequence. In our work, we demonstrate that intermediate estimates from diffusion models, for attributes like toxicity, can be used even with off-the-shelf reward models by evaluating the reward models on intermediate denoised estimates (see fig. 5).
@@ 4 Experiments
  • fk steering can improve gradient-guidance. In table 3, we show that fk steering SDv1.5 with $k=4$ outperforms gradient-guidance using the ImageReward model, using less sampling time. We note that using gradient-guidance for the proposal generator can improve performance even further, at the cost of increased sampling time and more compute.
@@
Model Sampler GenEval IR HPS Time
SDv1.5 $k=1$ 0.44 0.187 0.245 2.4s
SDv1.5 $\nabla$($k=1$) 0.45 0.668 0.245 20s
SDv1.5 fk($k=4$) 0.54 0.898 0.263 8.1s
SDv1.5 fk($\nabla$, $k=4$) 0.56 1.290 0.268 55s
Table 3: Comparison against gradient guidance. Here we note that fk steering with the model as the proposal generator outperforms gradient guidance, with faster sampling. We also note that fk steering can benefit from gradient guidance, albeit at the cost of more compute and sampling time.
@@
For fk steering, we resample $50$ times during the inference process (every $10$ steps for SSD-LM and every $20$ or MDLM). For all models we use $\lambda=10.0$, and return the highest scoring sample at $t=0$. Following Han et al. (2023), we generate $20$ continuations of length $50$ using their $15$ controllable generation prompts. In addition to fk steering, we evaluate base model performance and best-of-$n$. As a baseline to compare quality improvements from scaling inference steps (Gong et al., 2023; Sahoo et al., 2024), we also include results for SSD-LM with $T=5000$ (versus $T=500$). We additionally include results for GPT2-Medium (Radford et al., 2019b) as an auto-regressive baseline. We evaluate all methods using perplexity computed with GPT2-XL (Radford et al., 2019b), CoLA acceptability, and distinct uni/bi/trigrams (Dist-1/2/3) (Han et al., 2023; Li et al., 2016). Additional details on our text experiments are included in B.
@@
GPT2-medium $1$ 14.1 87.6 54/89/94
SSD-LM $1$ 23.2 68.3 46/83/92
SSD-LMT×10 $1$ 18.8 76.6 46/81/90
FK(GPT2) $4$ 11.0 80.0 40/73/86
FK(Trigram) $4$ 14.1 77.4 41/76/88
FK(CoLA) $4$ 17.4 95.7 45/80/90
BoN(GPT2) $4$ 13.6 75.6 42/77/88
BoN(Trigram) $4$ 15.9 71.9 43/78/89
BoN(CoLA) $4$ 19.2 93.8 46/81/91
BoN(GPT2) $8$ 11.2 80.3 41/74/86
BoN(Trigram) $8$ 13.9 76.8 42/77/89
BoN(CoLA) $8$ 18.4 97.2 46/82/91
MDLM $1$ 85.3 28.9 57/91/94
FK(GPT2) $4$ 49.0 39.8 52/86/92
FK(Trigram) $4$ 40.3 37.0 50/87/93
FK(CoLA) $4$ 73.6 69.8 59/88/91
BoN(GPT2) $4$ 55.5 32.9 54/88/93
BoN(Trigram) $4$ 52.1 30.1 53/89/93
BoN(CoLA) $4$ 71.4 59.4 57/90/94
BoN(GPT2) $8$ 46.9 37.2 53/87/92
BoN(Trigram) $8$ 45.9 35.4 52/88/93
BoN(CoLA) $8$ 68.2 73.1 58/91/94
Table 4: Text sample quality results metrics. We sample texts of length $50$ from all models and score perplexity with GPT2-XL and CoLA acceptability. Results are averaged over three seeds. Both SSD-LM and GPT-medium have $355$ million parameters. MDLM is a smaller model with $170$ million parameters.
@@ 4.3 Studying different choices of rewards in fk steering
SSD-LM 0.4% 1.2% 23.2 46/83/92
SSD-LM ($\nabla$ guidance) 22.3% 22.6% 40.3 53/89/94
MDLM 0.3% 1.9% 85.3 57/91/94
SSD-LM (no gradients)
BoN(4) 1.6% 4.8% 21.9 46/81/91
BoN(8) 5.0% 8.1% 23.0 46/81/91
FK($k=4$) 8.4% 14.0% 22.5 45/81/91
FK($k=4$, learned $r_{\phi}$) 15.2% 19.6% 26.3 45/83/91
FK($k=8$) 25.0% 29.7% 23.9 45/81/91
FK($k=8$, learned $r_{\phi}$) 39.0% 38.0% 26.9 45/83/91
MDLM (discrete, no gradients)
BoN(4) 2.2% 6.7% 83.8 57/90/93
BoN(8) 3.7% 10.8% 84.6 57/90/93
FK($k=4$) 23.0% 29.0% 81.0 56/90/93
FK($k=4$, many $r_{\phi}$) 37.0% 40.2% 83.0 57/90/92
FK($k=8$) 53.4% 48.3% 74.3 56/89/92
FK($k=8$, many $r_{\phi}$) 64.7% 51.7% 82.9 57/89/92
Table 6: Toxicity results. We evaluate the toxicity of the generated samples with (a) the classifier used for steering and (b) a separate holdout classifier, we also report GPT2-XL perplexity. Results are averaged over three seeds.
@@
We consider the task of red-teaming toxicity, a rare attribute identified in only $1\%$ of base SSD-LM samples and $0.3\%$ of MDLM samples. Toxicity presents an undesirable behavior for language models. Here, we examine whether fk steering enables examining rare but dangerous model behavior, a critical factor considered before deploying systems (Zhao et al., 2024a). In this experiment, we run fk steering with the base text diffusion models with SSD-LM (Han et al., 2023) and MDLM models. We use the base model as the proposal generator. As a baseline, we compare against gradient guidance for SSD-LM. For reward, we use a popular toxicity classifier (Logacheva et al., 2022). [^9] https://huggingface.co/s-nlp/roberta_toxicity_classifier
@@
  • Learned rewards improves controllability: With $8$ particles, fk steering SSD-LM outperforms gradient guidance on fraction of generations labeled toxic ($29.7\%$ vs $22.6\%$) by the holdout classifier, and dramatically improves on perplexity ($23.9$ vs $40.3$). Using improved intermediate rewards, learned from real data, improves this fraction even further to $39\%$.
  • Using many-sample $r_{\phi}$ improves controllability: fk steering MDLM with $k=8$ achieves an even higher fraction of text labeled toxic $48.3\%$. fk steering outperforms best-of-$n$ sampling with both $4,8$ particles. Using improved intermediate rewards for MDLM more samples for intermediate rewards improves performance even further to $51.7\%$ ($64.7\%$ on the guidance reward model).
@@
In the previous section, we observed that for the same number of particles, using either learned rewards or multiple samples improves performance. For instance, using $k=8$ with SSDLM without learning the intermediate rewards gets a toxicity rate of $25\%$, higher than the learned reward’s toxicity rate of $22.3\%$; using the learned rewards improves the toxicity rate to $39\%$. However, these rewards come with an added computational cost, either in training or more evaluations of the reward model, presenting a trade-off in terms of using extra computational resources either for more particles or for better intermediate rewards.
@@ 5 Conclusion
Promising future directions include exploring the value of varying the numbers of particles at inference time, either by having a variable budget or allocating compute adaptively. For applications like protein design, dynamically assigning greater numbers of particles to promising regions could enable generating a large number of diverse candidates. We also note that fk steering can also improve preference learning algorithms (Zhang and Ranganath, ) by improving the sample generator. Additionally, we believe it is crucial to better understand the limits of inference-time particle scaling and the corresponding compute-performance trade-offs.